# Output-Sensitive Evaluation of Regular Path Queries

### v1

@@ 6. When is OSPG better than PG?
Consider an input edge-labeled graph representing a path of length $N$ where all edges have a label ``$b"$. In particular, suppose that the vertices are $\{1,2,\ldots,N\}$ and the edges are $\{(1,``b",2),(2,``b",3),\ldots,(N-1,``b",N)\}$. Consider the RPQ $b^{*}c$. On this instance, the output is empty since there is no path with label $b^{*}c$. The NFA for $b^{*}c$ has a starting state $q_{0}$, a final state $q_{1}$, and transitions $(q_{0},``b",q_{0})$ and $(q_{0},``c",q_{1})$. The product graph has vertices $\{(1,q_{0}),\ldots,(N,q_{0})\}\cup\{(1,q_{1}),\ldots,(N,q_{1})\}$ and edges $\{((1,q_{0}),(2,q_{0})),\ldots,((N-1,q_{0}),(N,q_{0}))\}$.
  • The traditional PG algorithm takes time $\Omega(N^{2})$ on this instance. This is because for every one of the vertices $\{(1,q_{0}),\ldots,(N,q_{0})\}$ that correspond to the start state, it will run a BFS to discover a path to all subsequent vertices. Yet, it will never discover a way to extend any of these paths to reach any of the vertices $\{(1,q_{1}),\ldots,(N,q_{1})\}$ that correspond to the finish state.
@@
Figure 2 depicts this graph for $N=6$. Suppose that the RPQ is $ab^{*}c$. The output is empty. The NFA has states $\{q_{0},q_{1},q_{2}\}$ where $q_{0}$ is a start state, $q_{2}$ is a final state, and the transitions are $\{(q_{0},``a",q_{1}),(q_{1},``b",q_{1}),(q_{1},``c",q_{2})\}$. The product graph has edges $((v,q_{0}),(u,q_{1}))$ and $((v,q_{1}),$ $(u,q_{1}))$ for every pair of consecutive vertices $(v,u)$ in the first cycle. It also has edges $((v,q_{1}),(u,q_{1}))$ and $((v,q_{1}),(u,q_{2}))$ for every pair of consecutive vertices $(v,u)$ in the second cycle. The goal is to find a path from $(v,q_{0})$ to $(u,q_{2})$ for every pair of vertices $(v,u)$.
@@ 7.1. Kleene-free RPQs
The following theorem concerning the output-sensitive complexity of a $k$-path query was recently proved in (Hu, 2024):
Theorem 7.1 ($k$-path query (Hu, 2024)).
@@ 7.2. The RPQ $a^{*}$
8:     Evaluate using LeapFrogTrieJoin (Veldhuizen, 2014) or GenericJoin (Ngo et al., 2014) with variable order $(X,Z,Y)$:
@@
Consider Algorithm 2. Let $I$ be the last iteration, i.e. the smallest $i$ such that $\delta T^{i}=\emptyset$. Note that $T^{I}$ is the final output, i.e. $T^{I}\stackrel{{\scriptstyle\text{def}}}{{=}}T$. Moreover notice that by definition, the relations $\delta^{0}T,\ldots,\delta^{I}T$ form a partition of the output $T$. This is because Eq. (13) ensures that every $\delta^{i}T$ is disjoint from all the previous $\delta^{0}T,\ldots,\delta^{i-1}T$, thanks to the negation $\neg T^{i-1}(X,Y)$. Hence, the total time needed to evaluate the rule in Eq. (13) over all $i\in[I]$ is:
(15) $\sum_{i\in[I]}\sum_{(x,z)\in\delta T^{i-1}}|\sigma_{Z=z}E(Z,Y)|$ $\leq\sum_{(x,z)\in T}|\sigma_{Z=z}E(Z,Y)|$
@@
$=\sum_{(x,z)}|\sigma_{X=x,Z=z}T(X,Z)|^{1/2}\cdot|\sigma_{Z=z}E(Z,Y)|^{1/2}\cdot|\sigma_{X=x}T(X,Y)|^{1/2}$
@@
Inequality (15) holds because the relations $\delta^{0}T,\ldots,\delta^{I}T$ form a partition of the output $T$. Inequality (16) holds because by definition of transitive closure, for every tuple $(x,z,y)$ where $T(x,z)$ and $E(z,y)$ are true, $T(x,y)$ is also true. Inequality (17) holds by the query decomposition lemma, which we re-state below for completeness:
@@
$\sum_{i\in[I]}\sum_{(x,z)\in\delta T^{i-1}}|\sigma_{Z=z}E(Z,Y)|$ $\leq\sum_{(x,z)\in T}|\sigma_{Z=z}E(Z,Y)|$
$\leq\sum_{x\in\pi_{X}T}\sum_{z}|\sigma_{Z=z}E(Z,Y)|$
@@ 8. Related Work
To represent the set of possible paths matching a regular path query, a succinct and lossless representation has been introduced (Martens et al., 2023b) and recently connected to factorized representations (Olteanu and Závodný, 2015) of the output of conjunctive queries (Kimelfeld et al., 2023). The representation can be computed in linear time combined complexity and allows to perform operations like enumeration, counting, random sampling, grouping, and taking unions, which are common operations on the tabular representation of all paths. An in-depth analytical study has been made on SPARQL queries formulated by end users on top of graph-structured data in order to explain why regular path queries in graph database applications behave better than worst-case complexity results suggest (Bonifati et al., 2020; Martens and Trautner, 2019a).
@@
Our output-sensitive query evaluation algorithm OSPG follows a well-established line of work on output-sensitive algorithms, which we highlight next.
@@
A growing line of work decomposes the query processing task into a pre-processing phase, where a data structure representing succinctly the query output is constructed, and an enumeration phase, where each tuple in the query output is enumerated with some delay $d$. The overall data complexity is expressed as $O(N^{w}+d\cdot\textsf{OUT})$, where $N$ is the size of the input database, OUTis the output size, and $w$ is a width parameter that depends on the query structure. A large number of results can be explained in this complexity framework, as overviewed next.
@@
We first consider the case of $\alpha$-acyclic conjunctive queries. The classical Yannakakis algorithm, developed more than four decades ago, is the first output-sensitive evaluation algorithm for acylic queries: It runs in time $O(N+N\cdot\textsf{OUT})$ data complexity (Yannakakis, 1981), i.e., for $w=1$ and $d=O(N)$ (Bagan et al., 2007). For free-connex $\alpha$-acyclic conjunctive queries, $d$ decreases to $O(1)$ while $w$ remains 1 (Bagan et al., 2007). For hierarchical queries, which are a strict subset of the class of $\alpha$-acyclic conjunctive queries, $w=1+(\omega-1)\epsilon$ and $d=O(N^{1-\epsilon})$, where $\omega$ is the fractional hypertree width of the query (adapted from Boolean to conjunctive queries with arbitrary free variables) and $\epsilon\in[0,1]$ can be chosen arbitrarily (Kara et al., 2023c). There are similar trade-offs between $w$ and $d$ for general $\alpha$-acyclic queries (Kara et al., 2023b) and queries with access patterns (Kara et al., 2023a; Zhao et al., 2023). There are recent polynomial-time factor improvements on the Yannakakis algorithm (Deep et al., 2020; Hu, 2024, 2024; Deep et al., 2024), following an initial observation by Amossen and Pagh (Amossen and Pagh, 2009) on the output-sensitive evaluation of a two-relation join-project query that encodes the multiplication of two sparse matrices. Any $\alpha$-acyclic conjunctive query can be evaluated in time $O(N+\textsf{OUT}+N\cdot\textsf{OUT}^{5/6})$ using a non-combinatorial algorithm (using fast matrix multiplication) (Hu, 2024) and in time $O(N+\textsf{OUT}+N\cdot\textsf{OUT}^{1-\epsilon})$ using a combinatorial algorithm (Deep et al., 2024), where $\epsilon\in(0,1]$ is a query-dependent constant. These complexities also fit the general complexity framework by taking $w=1$ and $d=O(1+N/\textsf{OUT}^{1/6})$ and respectively $d=O(1+N/\textsf{OUT}^{1/k})$. The $k$-path query discussed in Section 7 can be evaluated in time $O(N+N\cdot\textsf{OUT}^{1/2})$ (Hu, 2024) or in time $O(N+\textsf{OUT}+N\cdot\textsf{OUT}^{1-1/k})$ (Deep et al., 2024). A lower bound of $\Omega(N+N\cdot\textsf{OUT}^{1/2})$ is also known for this query (Hu, 2024). This line of work culminates with an output-sensitive semi-ring extension of the Yannakakis algorithm to achieve $O(N+N\cdot\textsf{OUT}^{1-1/\textsf{outw}})$ data complexity [^8] We ignore here a $\text{\sf polylog}(N)$ factor. for any acyclic join-aggregate query, where outw is the so-called out-width of the query (Hu, 2024); this complexity fits the complexity framework with $w=1$ and $d=O(1+N/\textsf{OUT}^{1/\textsf{outw}})$. This upper bound is accompanied by a matching lower bound (modulo a $\text{\sf polylog}(N)$ factor).
@@
Arbitrary conjunctive queries (so including cyclic queries) can be evaluated using the factorized databases framework (Olteanu and Závodný, 2015) where the width $w$ is the fractional hypertree width [^9] This is generalized from Boolean to conjunctive queries; $w$ is denoted as $s$ in (Olteanu and Závodný, 2015) and FAQ-width in (Abo Khamis et al., 2016). and the delay $d=O(1)$. For conjunctive queries over the Boolean semiring, $w$ becomes the submodular width, which is smaller than the fractional hypertree width for many queries, and $d=O(\text{\sf polylog}(N))$ using the PANDA algorithm (Abo Khamis et al., 2017; Berkholz and Schweikardt, 2019). For conjunctive queries over arbitrary semirings, $w$ becomes the sharp submodular width, which is sandwiched between the fractional hypertree width and the submodular width, and $d$ remains $O(\text{\sf polylog}(N))$ using the FAQ-AI framework (Abo Khamis et al., 2020).
@@ Appendix A Transitive Closure: Linear versus Binary Formulation
In the proof of Theorem 7.4, we analyzed the overall runtime needed for semi-naïve evaluation of the linear formulation of transitive closure given by Eq. (10). By linear here, we mean that the recursive rule given by Eq. (10) has a single recursive atom $T(X,Z)$ in its body. In contrast, transitive closure can also be defined using a binary formulation, where the recursive rule has two recursive atoms in its body:
@@
We bound the total time needed to evaluate the rule in Eq. (22) over all iterations. Rule (23) is similar. Similar to the proof of Theorem 7.4, the relations $\delta T^{i}$ form a partition of the final $T$, thanks to the negation $\neg T^{i-1}(X,Y)$ in Eq. (22) and (23). Hence, the total time needed to evaluate Eq. (22) over all iterations is upper bounded by:
$\sum_{i\in[I]}\sum_{(x,z)\in\delta T^{i-1}}|\sigma_{Z=z}T(Z,Y)|$ $\leq\sum_{(x,z)\in T}|\sigma_{Z=z}T(Z,Y)|$
@@
$=\sum_{(x,z)}|\sigma_{X=x,Z=z}T(X,Z)|^{1/2}\cdot|\sigma_{Z=z}T(Z,Y)|^{1/2}\cdot|\sigma_{X=x}T(X,Y)|^{1/2}$
@@
Inequality (25) holds because by definition of transitive closure, for every tuple $(x,z,y)$ where $T(x,z)$ and $T(z,y)$ are true, $T(x,y)$ must be true as well. This proves that the total time needed to evaluate the rule in Eq. (22), and by extension Eq. (23), over all iterations is $O(\textsf{OUT}^{3/2})$. Rule (20) takes time $O(|E|)=O(\textsf{OUT})$, and rules (21) and (24) are dominated by other rules. This proves that the overall time for semi-naïve evaluation of the binary formulation of transitive closure is $O(\textsf{OUT}^{3/2})$.

### v2

@@ 6. When is OSPG strictly better than PG?
Consider an input edge-labeled graph representing a path of length $N$ where all edges have a label ``$b"$. In particular, suppose that the vertices are $\{1,2,\ldots,N\}$ and the edges are $\{(1,``b",2),(2,``b",3),\ldots,(N-1,``b",N)\}$. Consider the RPQ $b^{*}c$. On this instance, the output is empty since there is no path with label $b^{*}c$. The NFA for $b^{*}c$ has an initial state $q_{0}$, an accepting state $q_{1}$, and transitions $(q_{0},``b",q_{0})$ and $(q_{0},``c",q_{1})$. The product graph has vertices $\{(1,q_{0}),\ldots,(N,q_{0})\}\cup\{(1,q_{1}),\ldots,(N,q_{1})\}$ and edges $\{((1,q_{0}),(2,q_{0})),\ldots,((N-1,q_{0}),(N,q_{0}))\}$.
  • The traditional PG algorithm takes time $\Omega(N^{2})$ on this instance. This is because for every one of the vertices $\{(1,q_{0}),\ldots,(N,q_{0})\}$ that correspond to the initial state, it will run a BFS to discover a path to all subsequent vertices. Yet, it will never discover a way to extend any of these paths to reach any of the vertices $\{(1,q_{1}),\ldots,(N,q_{1})\}$ that correspond to the accepting state.
@@
Figure 2 depicts this graph for $N=6$. Suppose that the RPQ is $ab^{*}c$. The output is empty. The NFA has states $\{q_{0},q_{1},q_{2}\}$ where $q_{0}$ is an initial state, $q_{2}$ is an accepting state, and the transitions are $\{(q_{0},``a",q_{1}),(q_{1},``b",q_{1}),(q_{1},``c",q_{2})\}$. The product graph has edges $((v,q_{0}),(u,q_{1}))$ and $((v,q_{1}),$ $(u,q_{1}))$ for every pair of consecutive vertices $(v,u)$ in the first cycle. It also has edges $((v,q_{1}),(u,q_{1}))$ and $((v,q_{1}),(u,q_{2}))$ for every pair of consecutive vertices $(v,u)$ in the second cycle. The goal is to find a path from $(v,q_{0})$ to $(u,q_{2})$ for every pair of vertices $(v,u)$.
@@ 7.1. Kleene-free RPQs
The following theorem concerning the output-sensitive complexity of a $k$-path query was recently proved in (Hu, 2025):
Theorem 7.1 ($k$-path query (Hu, 2025)).
@@ 7.2. The RPQ $a^{*}$
8:     Evaluate $\delta T^{i}$ below by iterating over $(x,z)\in\delta T^{i-1}$, then over neighbors $y$ of $z$, then checking that $\neg T^{i-1}(x,y)$ holds, and if so, adding $(x,y)$ to $\delta T^{i}$:
@@
Consider Algorithm 2. Let $I$ be the last iteration, i.e. the smallest $i$ such that $\delta T^{i}=\emptyset$. Note that $T^{I}$ is the final output, i.e. $T^{I}\stackrel{{\scriptstyle\text{def}}}{{=}}T$. Moreover notice that by definition, the relations $\delta^{0}T,\ldots,\delta^{I}T$ form a partition of the output $T$. This is because Eq. (13) ensures that every $\delta^{i}T$ is disjoint from all the previous $\delta^{0}T,\ldots,\delta^{i-1}T$, thanks to the negation $\neg T^{i-1}(X,Y)$. For a fixed $i$, we evaluate the rule in Eq. (13) by going through all tuples $(x,z)$ in $\delta^{i-1}T$, and for each one of them, we go over all neighbors $y$ of $z$, then check whether the negation $\neg T^{i-1}(x,y)$ holds, and if so, we add $(x,y)$ to $\delta T^{i}$.  [^8] This evaluation strategy is equivalent to using LeapFrogTrieJoin (Veldhuizen, 2014) or GenericJoin (Ngo et al., 2014) to compute the join $\delta T^{i-1}(X,Z)\land E(Z,Y)$ with the variable order $(X,Z,Y)$, and then filtering the result using the negation $\neg T^{i-1}(X,Y)$. Therefore, the total time needed to evaluate the rule in Eq. (13) over all iterations $i\in[I]$ is:
(15) $\sum_{i\in[I]}\sum_{(x,z)\in\delta T^{i-1}}|\sigma_{Z=z}E(Z,Y)|$ $=\sum_{(x,z)\in T}|\sigma_{Z=z}E(Z,Y)|$
$=\sum_{(x,z)\in T}|\sigma_{Z=z}E(Z,Y)|^{1/2}\cdot|\sigma_{Z=z}E(Z,Y)|^{1/2}$
@@
$=\sum_{(x,z)\in V^{2}}|\sigma_{X=x,Z=z}T(X,Z)|^{1/2}\cdot|\sigma_{Z=z}E(Z,Y)|^{1/2}\cdot|\sigma_{X=x}T(X,Y)|^{1/2}$
@@
Equality (15) holds because the relations $\delta^{0}T,\ldots,\delta^{I}T$ form a partition of the output $T$, and $\delta^{I}T=\emptyset$. Inequality (16) holds because by definition of transitive closure, for every tuple $(x,z,y)$ where $T(x,z)$ and $E(z,y)$ are true, $T(x,y)$ is also true. This implies that for every $(x,z)\in T$, we have $\pi_{Y}\left(\sigma_{Z=z}E(Z,Y)\right)\subseteq\pi_{Y}\left(\sigma_{X=x}T(X,Y)\right)$, hence $|\sigma_{Z=z}E(Z,Y)|\leq|\sigma_{X=x}T(X,Y)|$. Inequality (17) holds by the query decomposition lemma, which we re-state below for completeness:
@@
$\sum_{i\in[I]}\sum_{(x,z)\in\delta T^{i-1}}|\sigma_{Z=z}E(Z,Y)|$ $=\sum_{(x,z)\in T}|\sigma_{Z=z}E(Z,Y)|$
$\leq\sum_{x\in\pi_{X}T}\sum_{z\in V}|\sigma_{Z=z}E(Z,Y)|$
@@ 8. Related Work
To represent the set of possible paths matching a regular path query, a succinct and lossless representation has been introduced (Martens et al., 2023b) and recently connected (Kimelfeld et al., 2025) to factorized representations (Olteanu and Závodný, 2015) of the output of conjunctive queries. The representation can be computed in linear time combined complexity and allows to perform operations like enumeration, counting, random sampling, grouping, and taking unions, which are common operations on the tabular representation of all paths. An in-depth analytical study has been made on SPARQL queries formulated by end users on top of graph-structured data in order to explain why regular path queries in graph database applications behave better than worst-case complexity results suggest (Bonifati et al., 2020; Martens and Trautner, 2019a).
@@
Our output-sensitive query evaluation algorithm OSPG follows a well-established line of work on output-sensitive algorithms, which we overview next.
@@
A growing line of work decomposes the query processing task into a pre-processing phase, where a data structure representing succinctly the query output is constructed, and an enumeration phase, where each tuple in the query output is enumerated with some delay $d$. The overall data complexity is expressed as $O(N^{w}+d\cdot\textsf{OUT})$, where $N$ is the size of the input database, OUT is the output size, and $w$ is a width parameter that depends on the query structure. A large number of results can be explained in this complexity framework, as exemplified next.
@@
We first consider the case of $\alpha$-acyclic conjunctive queries. The classical Yannakakis algorithm, developed more than four decades ago, is the first output-sensitive evaluation algorithm for acyclic queries: It runs in time $O(N+N\cdot\textsf{OUT})$ data complexity (Yannakakis, 1981), i.e., for $w=1$ and $d=O(N)$ (Bagan et al., 2007). For free-connex $\alpha$-acyclic conjunctive queries, $d$ decreases to $O(1)$ while $w$ remains 1 (Bagan et al., 2007). For hierarchical queries, which are a strict subset of the class of $\alpha$-acyclic conjunctive queries, $w=1+(\omega-1)\epsilon$ and $d=O(N^{1-\epsilon})$, where $\omega$ is the fractional hypertree width of the query (adapted from Boolean to conjunctive queries with arbitrary free variables) and $\epsilon\in[0,1]$ can be chosen arbitrarily (Kara et al., 2023c). There are similar trade-offs between $w$ and $d$ for general $\alpha$-acyclic queries (Kara et al., 2023b) and queries with access patterns (Kara et al., 2023a; Zhao et al., 2023). There are recent polynomial-time factor improvements on the Yannakakis algorithm (Deep et al., 2020; Hu, 2024, 2025; Deep et al., 2024), following an initial observation by Amossen and Pagh (Amossen and Pagh, 2009) on the output-sensitive evaluation of a two-relation join-project query that encodes the multiplication of two sparse matrices. Any $\alpha$-acyclic conjunctive query can be evaluated in time $O(N+\textsf{OUT}+N\cdot\textsf{OUT}^{5/6})$ using a non-combinatorial algorithm (using fast matrix multiplication) (Hu, 2024) and in time $O(N+\textsf{OUT}+N\cdot\textsf{OUT}^{1-\epsilon})$ using a combinatorial algorithm (Deep et al., 2024), where $\epsilon\in(0,1]$ is a query-dependent constant. These complexities also fit the general complexity framework by taking $w=1$ and $d=O(1+N/\textsf{OUT}^{1/6})$ and respectively $d=O(1+N/\textsf{OUT}^{1/k})$. The $k$-path query discussed in Section 7 can be evaluated in time $O(N+N\cdot\textsf{OUT}^{1/2})$ (Hu, 2025) or in time $O(N+\textsf{OUT}+N\cdot\textsf{OUT}^{1-1/k})$ (Deep et al., 2024). A lower bound of $\Omega(N+N\cdot\textsf{OUT}^{1/2})$ is also known for this query (Hu, 2025). This line of work culminates with an output-sensitive semi-ring extension of the Yannakakis algorithm to achieve $O(N+\textsf{OUT}+N\cdot\textsf{OUT}^{1-1/\textsf{outw}})$ data complexity [^9] We ignore here a $\text{\sf polylog}(N)$ factor. for any acyclic join-aggregate query, where outw is the so-called out-width of the query (Hu, 2025); this complexity fits the complexity framework with $w=1$ and $d=O(1+N/\textsf{OUT}^{1/\textsf{outw}})$. This upper bound is accompanied by a matching lower bound (modulo a $\text{\sf polylog}(N)$ factor).
@@
Arbitrary conjunctive queries (so including cyclic queries) can be evaluated using the factorized databases framework (Olteanu and Závodný, 2015) where the width $w$ is the fractional hypertree width [^10] This is generalized from Boolean to conjunctive queries; $w$ is denoted as $s^{\uparrow}$ in (Olteanu and Závodný, 2015) and FAQ-width in (Abo Khamis et al., 2016). and the delay $d=O(1)$. For conjunctive queries over the Boolean semiring, $w$ becomes the submodular width, which is smaller than the fractional hypertree width for many queries, and $d=O(\text{\sf polylog}(N))$ using the PANDA algorithm (Abo Khamis et al., 2017; Berkholz and Schweikardt, 2019). For conjunctive queries over arbitrary semirings, $w$ becomes the sharp submodular width, which is sandwiched between the fractional hypertree width and the submodular width, and $d$ remains $O(\text{\sf polylog}(N))$ using the FAQ-AI framework (Abo Khamis et al., 2020).
@@
Output-sensitive transitive closure.
Let $G=(V,E)$ be a directed graph. Using fast matrix multiplication, it is possible to compute the transitive closure of $G$ in time $\tilde{O}(|V|^{\omega})$ (Kozen, 1992), where $\omega$ is the matrix multiplication exponent, i.e. the smallest exponent $\alpha$ that allows us to multiply two $N\times N$ matrices in time $\tilde{O}(N^{\alpha})$, and $\tilde{O}$ hides a factor of $\text{\sf polylog}(N)$. Currently, the best known upper bound for $\omega$ is $2.371552$ (Vassilevska Williams et al., 2024). An output-sensitive extension of this result computes the transitive closure of $G$ in time $\tilde{O}(|V|^{\frac{\omega+1}{4}}\cdot\textsf{OUT})$ (Borassi et al., 2015), which translates to $\tilde{O}(|V|^{0.842888}\cdot\textsf{OUT})$ using the best known $\omega$, where OUT is the number of edges in the transitive closure. Without using fast matrix multiplication, (Gucht et al., 2015) presents a combinatorial algorithm for transitive closure that runs in time $O(\textsf{OUT}^{3/2})$. Using fast matrix multiplication, (Abboud et al., 2024) improves this bound to $\tilde{O}(\textsf{OUT}^{1.3459})$, under the best known $\omega$. Note that in transitive closure, the input graph is always a subset of the output, hence $|E|\leq\textsf{OUT}$. Therefore, our OSPG transitive closure bound, $O(\textsf{OUT}\cdot\sqrt{|E|})$, is never larger than the bound of $O(\textsf{OUT}^{3/2})$ from (Gucht et al., 2015), and is strictly smaller whenever $\textsf{OUT}>|E|$. Moreover, our OSPG bound is strictly smaller than the bound of $O(\textsf{OUT}^{1.3459})$ from (Abboud et al., 2024) whenever $\textsf{OUT}>|E|^{1.4455}$.
@@ Appendix A Transitive Closure: Linear versus Binary Formulation
In the proof of Theorem 7.4, we analyzed the overall runtime needed for semi-naïve evaluation of the linear formulation of transitive closure given by Eq. (10). By linear here, we mean that the recursive rule given by Eq. (10) has a single recursive atom $T(X,Z)$ in its body. In contrast, the transitive closure can also be defined using a binary formulation, where the recursive rule has two recursive atoms in its body:
@@
Let $I$ be the last iteration, i.e. the smallest $i$ such that $\delta T^{i}=\emptyset$. We bound the total time needed to evaluate the rule in Eq. (22) over all iterations. Rule (23) is similar. Similar to the proof of Theorem 7.4, the relations $\delta T^{i}$ form a partition of the final $T$, thanks to the negation $\neg T^{i-1}(X,Y)$ in Eq. (22) and (23). Hence, the total time needed to evaluate Eq. (22) over all iterations is upper bounded by:
$\sum_{i\in[I]}\sum_{(x,z)\in\delta T^{i-1}}|\sigma_{Z=z}T(Z,Y)|$ $=\sum_{(x,z)\in T}|\sigma_{Z=z}T(Z,Y)|$
$=\sum_{(x,z)\in T}|\sigma_{Z=z}T(Z,Y)|^{1/2}\cdot|\sigma_{Z=z}T(Z,Y)|^{1/2}$
@@
$=\sum_{(x,z)\in V^{2}}|\sigma_{X=x,Z=z}T(X,Z)|^{1/2}\cdot|\sigma_{Z=z}T(Z,Y)|^{1/2}\cdot|\sigma_{X=x}T(X,Y)|^{1/2}$
@@
Inequality (25) holds because by definition of transitive closure, for every tuple $(x,z,y)$ where $T(x,z)$ and $T(z,y)$ are true, $T(x,y)$ must be true as well. This proves that the total time needed to evaluate the rule in Eq. (22), and by symmetry Eq. (23), over all iterations is $O(\textsf{OUT}^{3/2})$. Rule (20) takes time $O(|E|)=O(\textsf{OUT})$, and rules (21) and (24) are dominated by other rules. This proves that the overall time for semi-naïve evaluation of the binary formulation of transitive closure is $O(\textsf{OUT}^{3/2})$.